Jose takes twice as long as Maria to sort the mail. Together they can sort the mail in 8 hours. How long does it take each of them to sort the mail?
Answer by richwmiller(17219)

8/x+8/2x=1
8/x+4/x=1
12/x=1
12=x maria
24=2x jose
check 8/12+8/24=1
2/3+1/3=1
ok